# Does a Rising Tide Lift All Boats? Bias Mitigation for AI-based CMR Segmentation

### v1

@@ Abstract
Artificial intelligence (AI) is increasingly being used for medical imaging tasks. However, there can be biases in the resulting models, particularly when they were trained using imbalanced training datasets. One such example has been the strong race bias effect in cardiac magnetic resonance (CMR) image segmentation models. Although this phenomenon has been reported in a number of publications, little is known about the effectiveness of bias mitigation algorithms in this domain. We aim to investigate the impact of common bias mitigation methods to address bias between Black and White subjects in AI-based CMR segmentation models. Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the race bias. Furthermore, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images. We find that bias can be mitigated using oversampling, significantly improving performance for the underrepresented Black subjects whilst not significantly reducing the majority White subjects’ performance. Group DRO also improves performance for Black subjects but not significantly, while reweighing decreases performance for Black subjects. Using a combination of oversampling and Group DRO also improves performance for Black subjects but not significantly. Using cropped images increases performance for both races and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further.
Keywords:
Bias mitigation CMR imaging Segmentation.
@@ 1 Introduction
Artificial intelligence (AI) is increasingly being used to aid medical diagnosis, prognosis and treatment planning. However, AI models have been shown to exhibit bias by protected attributes in many different applications. For example, in cardiac magnetic resonance (CMR) image segmentation, both sex [1] and race [2, 3, 4] bias have been reported. Sex bias has also been found in chest X-ray image classification [5, 6]. AI bias can have detrimental downstream impacts in medical imaging applications. For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes [4].
Previous work has aimed to address biases in AI models for medical imaging tasks by using bias mitigation methods. [7] evaluated and mitigated bias in chest X-ray classification models using methods based on oversampling minority protected groups and Pareto minimax optimisation. [8] proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. [9] evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation. In CMR segmentation, [3] proposed oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
@@
Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias [10]. One key finding of this work was that distributional differences in areas outside the heart were significant in the learning of biased representations in AI segmentation models. This finding opens up new avenues of enquiry in bias mitigation, which we explore in this paper.
The contribution of this paper is to perform the most extensive and comprehensive investigation to date of bias mitigation methods in AI-based CMR segmentation. We investigate multiple mitigation methods as well as combinations of these methods. We also investigate AI models trained using CMR images that are cropped to remove features outside of the heart that may be leading to bias, and evaluate the effectiveness of bias mitigation techniques in this setting. Our results show that CMR segmentation bias can be mitigated to produce results that have high accuracy but are not significantly different between races.
2 Methods
@@
We use CMR images from the UK Biobank [11]. The dataset consists of end diastolic (ED) and end systolic (ES) cine short-axis images from 5,778 subjects. Manual segmentation of the left ventricular blood pool (LVBP), left ventricular myocardium (LVM), and right ventricular blood pool (RVBP) was performed for the ED and ES images of each subject. The LV endocardial and epicardial borders and the RV endocardial border were outlined using cvi42 (version 5.1.1, Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada). The same guidelines were provided to a panel of ten experts with one expert annotating each image. Each expert was provided with a random selection of images for annotation which included subjects of different sexes and races. They were not provided with demographic information about the subjects.
Previous work [2, 1] has shown that bias is greater when the imbalance between races in the training set is greater. Therefore, we curate a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our training set comprises 15 Black subjects and 4,221 White subjects.
@@ 2.1 Baseline Model
We trained a baseline 2D nnU-Net v1 model [12] to segment the LVM, LVBP and RVBP using the training parameters from [1]. The demographic information of the subjects used in the training and testing datasets can be seen in Table 1 and Table 2. Training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other.
Health measure Overall White Black # subjects 4236 4221 15 Age (years 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Standing height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)*
Table 2: Characteristics of subjects used in the test dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Mixed Asian Black Chinese Other # subjects 1542 469 170 387 223 111 182 Age 61.7 (7.9) 64.8 (7.7)* 59.8 (7.2)* 61.0 (8.2)* 59.1 (7.1)* 59.7 (6.4)* 62.1 (7.5) Standing height 167.4 (9.2) 169.8 (9.4)* 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1)* 165.4 (9.7)* Weight 75.1 (15.1) 77.7 (15.1)* 75.1 (15.3) 72.3 (12.8)* 82.0 (16.0)* 63.5 (9.9)* 73.3 (15.5) Body Mass Index 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7)* 28.8 (5.1)* 24.0 (3.1)* 26.6 (4.4)
@@ 2.2 Oversampling
We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection [13]. Oversampling refers to the process of increasing the sampling of a minority group in a dataset. Here, we oversample the Black subjects in the training dataset so that they were equal to the number of White subjects in each batch used during training. This was performed using random sampling with replacement so each subject could in principle be selected more than once in a training batch. The same training parameters from [1] were used.
@@ 2.3 Reweighing
An nnU-Net model was also trained using a reweighing mitigation strategy [13]. Reweighing refers to the process of increasing the importance of under-represented groups to the model. This can be done as a pre-processing step or during training by adapting the loss function. We performed this during training by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net. Each group was weighted inversely proportionally to the group size, as shown in Eq. 2. The weights were then normalised so that they summed to 1. The training parameters from [1] were again used.
@@ 2.4 Group Distributionally Robust Optimisation
The final mitigation approach was Group Distributionally Robust Optimisation, or Group DRO, which was first proposed in [14]. The method aims to optimise the performance of the worst-performing group in a dataset. The Group DRO loss function can be formalised as:
@@
In this method, CE loss was used instead of CE-Dice loss which was used for the oversampling and reweighing experiments. Group DRO uses losses from individual samples to calculate the average loss for the groups. However, Dice loss is calculated using global statistics of the true positives, false positives and false negatives for a group or batch. It is non-additive as the numerator and denominator will change if the calculation is performed on a per-sample basis rather than for a group or batch. For Dice loss, the average group loss and global loss are different, which causes instability in training.
The code for the bias mitigation models is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation
@@ 2.5 Training Using Cropped Images
Following the findings of [10], which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images. The images were cropped around the heart using a bounding box defined based on the ground truth segmentation. All images were cropped to the same size, i.e. the size of the largest heart in the dataset.
2.6 Training Using Combinations of Mitigation Methods
The final experiment combined the mitigation methods into pairs to test whether combinations of methods would improve performance. This results in three additional methods: oversampling + Group DRO, reweighing + Group DRO, and oversampling + reweighing. These methods were applied in the same way as above. CE loss was used for experiments with Group DRO and CE-Dice loss was used for oversampling + reweighing.
2.7 Metrics
For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) for subjects in the test set. Performance was measured using the median and inter-quartile range (IQR) of the DSC. Fairness metrics are also reported: the fairness gap (FG) as defined in [1] and skewed error ratio (SER) as defined in [3]. The SER is defined as $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$. This measures the ratio of the errors between the median DSCs for the worst-performaing and best-performing protected groups. A more fair model will have a FG closer to 0 and a SER closer to 1.

### v2

@@ Abstract
Artificial intelligence (AI) is increasingly being used for medical imaging tasks. However, there can be biases in AI models, particularly when they are trained using imbalanced training datasets. One such example has been the strong ethnicity bias effect in cardiac magnetic resonance (CMR) image segmentation models. Although this phenomenon has been reported in a number of publications, little is known about the effectiveness of bias mitigation algorithms in this domain. We aim to investigate the impact of common bias mitigation methods to address bias between Black and White subjects in AI-based CMR segmentation models. Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the ethnicity bias. Second, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images. We find that bias can be mitigated using oversampling, significantly improving performance for the underrepresented Black subjects whilst not significantly reducing the majority White subjects’ performance. Using cropped images increases performance for both ethnicities and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further. When testing the models on an external clinical validation set, we find high segmentation performance and no statistically significant bias.
@@ 1 Introduction
Artificial intelligence (AI) is increasingly being used to aid medical diagnosis, prognosis and treatment planning. However, AI models have been shown to exhibit bias by protected attributes in many different applications (Larrazabal et al., 2020; Seyyed-Kalantari et al., 2021a; Klingenberg et al., 2023), including AI-based segmentation of cardiac magnetic resonance (CMR) images (Puyol-Antón et al., 2021, 2022; Lee et al., 2022, 2023). AI bias can have detrimental downstream impacts in medical imaging applications. For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes (Puyol-Antón et al., 2022).
Previous work has aimed to address biases in AI models for medical imaging tasks by using generic bias mitigation methods. For example, Zong et al. (2022) proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ Li and Li (2024). Furthermore, research has suggested that, whilst such generic bias mitigation approaches may appear to reduce bias when evaluated on internal validation sets, the fairness gains often do not hold when evaluated externally (Schrouff et al., 2022; Yang et al., 2024). We hypothesise that a possible reason for this lack of effectiveness is that these methods take a “blind” approach to addressing bias in AI models, i.e. they apply a generic bias mitigation algorithm that does not take account of the underlying cause of the bias. More recently, a new strand of research has emerged which attempts to understand the bias, with a view to using this knowledge to develop more informed mitigation approaches. For example, Glocker et al. (2023) analysed bias in chest X-ray classifiers, applying test-set resampling, multitask learning, and model inspection to provide insights into the way protected attributes were encoded in the AI model. Similarly, Olesen et al. (2025) used ‘slice discovery methods’ to reveal the root causes of sex bias, also in chest X-ray classifiers. Such methods potentially enable the development of bias mitigation techniques that are targeted at the underlying cause(s) of the bias, and could lead to improved and more robust mitigation, which will hold under the effects of domain shift to external validation sets.
@@
Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias (Lee et al., 2025a). One key finding of this work was that distributional differences in areas outside the heart were significant in the learning of biased representations in AI segmentation models. This finding opens up new avenues of enquiry in bias mitigation, which we explore in this paper.
2 Related Works
2.1 Bias in medical imaging
AI bias has been observed in a wide range of medical imaging modalities and tasks. For example, Seyyed-Kalantari et al. (2021a) found bias in chest X-ray classification models in terms of sex and ethnicity. A CNN-based model showed a favourable bias towards males and older people, with these groups having higher true positive rates. Similarly, Seyyed-Kalantari et al. (2021b) found that younger people, females, patients under 20 years old, Black patients and Hispanic patients had higher rates of under-diagnosis. Larrazabal et al. (2020) also used chest X-ray data, and investigated the effect of training a thoracic disease classification model on datasets that were imbalanced by sex. They found a relationship between training set representation and performance for both sexes, and that training a model with more balanced data did not significantly decrease accuracy for the majority group but significantly improved performance for the under-represented group.
Bias in brain magnetic resonance imaging (MRI) classification has also been reported. A key early work here was Petersen et al. (2022), who reported bias in a CNN-based model for Alzheimer’s disease (AD) detection. Similarly, Klingenberg et al. (2023) found higher accuracy in MRI-based AD detection in females than males, despite balancing the training dataset by age and sex. Wang et al. (2023) also assessed bias in brain MRI disease classification, investigating the impact of training choices on bias.
In dermatology imaging , Abbasi-Sureshjani et al. (2020) found bias by age and sex when training lesion classification models. Daneshjou et al. (2022) found disparities in accuracy between lighter and darker skin tones, with the model achieving higher accuracy on the lighter skin tones. Fine-tuning the model on a diverse dataset closed this gap in performance and improved overall model performance so that it was equal to, or exceeded, clinician accuracy.
Puyol-Antón et al. (2021) was the first paper to report bias in AI-based segmentation models. This work found that an nnU-Net model trained on ethnicity-imbalanced CMR data produced biased results by ethnicity. The segmentation performance favoured White subjects, who were in the majority in the training set. Subsequent work has investigated the downstream clinical impact of this bias (Puyol-Antón et al., 2022) and analysed it in more controlled experiments by both ethnicity and sex (Lee et al., 2022, 2023). Segmentation bias has since also been reported in brain MRI (Ioannou et al., 2022), dermatology images (Benčević et al., 2024) and orthopaedic radiographs (Siddiqui et al., 2024).
2.2 Bias mitigation methods
Bias mitigation methods aim to reduce the bias observed between protected groups in an AI model. They can be applied as pre-, in- or post-processing methods (Mehrabi et al., 2019). Pre-processing methods aim to transform the data before training to remove bias. These methods include targeted data collection and importance reweighing (which can also be performed as an in-processing method). Efforts to collect data from more diverse populations include a dataset of dermatology data from four African countries (Gottfrois et al., 2024), a leprosy skin imaging dataset from Brazil (Barbieri et al., 2022), brain MRI images from women with fibromyalgia in Mexico (Balducci et al., 2022) and brain tumour segmentation data from Nigeria (Adewole et al., 2024).
In-processing methods can be used to change the objective function or apply constraints to the model during training to reduce bias. An example of such an approach is Group Distributionally Robust Optimisation, or Group DRO (Sagawa et al., 2020), which alters the loss function to optimise performance for the worst performing group in a dataset. Group DRO was one of the approaches evaluated for mitigating bias in chest X-ray classifiers in the comparative analysis by Zhang et al. (2022). Another example of a modified loss function is Pareto minimax optimisation (Martinez et al., 2020), which aims to minimise an importance-weighted maximum ‘risk’ across protected groups. Adversarial learning has also been suggested as an in-processing method for bias mitigation. In Madras et al. (2018), a model was adversarially trained to learn fair representations using an encoder-decoder network. Similarly, Zhang et al. (2018) used adversarial debiasing with the added constraint of satisfying fairness definitions such as demographic parity, equalised odds or equalised opportunity.
Post-processing methods aim to modify the predictions made by the model based on a subject’s protected attributes. Lohia et al. (2019) proposed an algorithm which assesses an individual sample’s prediction, establishes whether it is biased and changes the prediction to the privileged group’s label if the sample experiences bias. Reject option classification, proposed in Kamiran et al. (2012), considers the confidence of predictions. For a binary classifier, prediction probabilities close to 0 or 1 represent confident predictions, whereas probabilities close to 0.5 represent more uncertain predictions. Samples are not assigned a label if their probabilities lie within a certain uncertainty range as they are considered more prone to bias and are relabelled depending on the group they belong to.
Bias mitigation in AI-based segmentation has received relatively little attention compared to classification tasks. Relevant works here include FairSeg (Tian et al., 2023), which published a fairness dataset for medical image segmentation and proposed a fair error-bound scaling approach to reweight the loss function. Siddiqui et al. (2024) evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation. Finally, in CMR segmentation, Puyol-Antón et al. (2021) internally evaluated protected group-specific models, oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
3 Contributions
In the context of AI-based segmentation of cine CMR, the main contributions of this novel work are:
  1. 1.
    We perform the most extensive and comprehensive investigation to date of multiple bias mitigation methods in CMR segmentation, as well as combinations of these methods.
  2. 2.
    We perform one of the first investigations into using knowledge of the root cause of bias for mitigation. Specifically, we train AI segmentation models using CMR images that are cropped to remove features outside of the heart. We evaluate a baseline model and state-of-the-art bias mitigation techniques in this setting.
  3. 3.
    We evaluate the efficacy of all approaches under both internal and external validation settings.
A preliminary version of this work has been published in Lee et al. (2025b). This paper extends that work by (i) incorporating a wider range of metrics and more in-depth analysis and discussion, (ii) inclusion of a new clinically-applicable ‘cascaded’ cropping based mitigation approach and (iii) including external validation of all approaches on a clinical dataset.
4 Materials and Methods
4.1 Data
@@
To train and internally validate all models we used CMR images from the UK Biobank (Petersen et al., 2016). The dataset consists of end diastolic (ED) and end systolic (ES) cine short-axis images from 5,778 subjects. Manual segmentation of the left ventricular blood pool (LVBP), left ventricular myocardium (LVM), and right ventricular blood pool (RVBP) was performed for the ED and ES images of each subject. The LV endocardial and epicardial borders and the RV endocardial border were outlined using cvi42 (version 5.1.1, Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada). The same guidelines were provided to a panel of ten experts with one expert annotating each image. Each expert was provided with a random selection of images for annotation which included subjects of different sexes and ethnicities. They were not provided with demographic information about the subjects.
Previous work (Lee et al., 2022, 2023) has shown that bias is greater when the imbalance between ethnicities in the training set is greater. Therefore, we curated a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset. The remaining subjects were used as the internal validation test set. The demographic information of the subjects in the training and test sets can be seen in Table 1 and Table 2. Note that training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other. All ethnicity information refers to self-reported ethnicity based on the categories provided in the UK Biobank dataset.
In addition, for external validation a dataset of cine short-axis CMR images from St. Thomas’ Hospital, London was used. All subjects were scanned using a 1.5T MRI scanner (Aera-Magneton, Siemens Healthcare, Erlangen, Germany) between November 2019 and November 2021. The patients scanned had suspected hypertension. Ethnicity was recorded as Black if both parents self-identified as African descendants or White if both parents self-identified as European descendants. The demographic information for the subjects used can be seen in Table 3. Ground truth segmentations were performed manually by clinical experts using cvi42 (Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada). Further details of the imaging protocol can be found in Georgiopoulos et al. (2024).
Health measure Overall White Black n subjects 4236 4221 15 Age (years) 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)
Table 2: Characteristics of subjects used in the internal validation test dataset. Mean (standard deviation) values are presented for each characteristic. One-way ANOVA was used to test for differences across ethnic groups. Statistically significant differences between groups are indicated with an asterisk. * (p $<$ 0.01).
Health measure Overall White Mixed Asian Black Chinese Other n subjects 1542 469 170 387 223 111 182 Age (years) * 61.7 (7.9) 64.8 (7.7) 59.8 (7.2) 61.0 (8.2) 59.1 (7.1) 59.7 (6.4) 62.1 (7.5) Height (cm) * 167.4 (9.2) 169.8 (9.4) 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1) 165.4 (9.7) Weight (kg) * 75.1 (15.1) 77.7 (15.1) 75.1 (15.3) 72.3 (12.8) 82.0 (16.0) 63.5 (9.9) 73.3 (15.5) Body Mass Index * 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7) 28.8 (5.1) 24.0 (3.1) 26.6 (4.4)
Table 3: Characteristics of subjects used in the external validation test set. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Black n subjects 84 30 54 Age (years) 46.0 (12.8) 40.5 (14.1) 49.1 (11.0) Weight (kg) 94.6 (19.7) 92.6 (19.3) 95.1 (20.2) Height (cm) 174.3 (8.96) 176.9 (7.33) 172.9 (9.48) Body Mass Index 31.2 (6.23) 29.6 (6.13) 31.8 (6.23)
@@ 4.2 Baseline Model
As a baseline model, we trained a 2D nnU-Net v1 model (Isensee et al., 2020) using the UK Biobank training set to segment the LVM, LVBP and RVBP.
@@ 4.3 Oversampling
We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection (Kamiran and Calders, 2012). Oversampling refers to the process of increasing the sampling of a minority group in a dataset. Here, we oversample the Black subjects in the training set so that they were equal to the number of White subjects in each batch used during training. This was performed using random sampling with replacement so each subject could in principle be selected more than once in a training batch.
@@ 4.4 Reweighing
An nnU-Net model was also trained using a reweighing mitigation strategy (Kamiran and Calders, 2012). Reweighing refers to the process of increasing the importance of under-represented groups to the model. We implemented this strategy by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net. Each group was weighted inversely proportionally to the group size, as shown in Eq. 1. The weights were then normalised so that they summed to 1, as shown in Eq. 2.
@@ 4.5 Group Distributionally Robust Optimisation
The final mitigation approach was Group DRO, which was first proposed in Sagawa et al. (2020). The method aims to optimise the performance of the worst-performing group in a dataset. The Group DRO loss function can be formalised as:
@@
In this method, CE loss was used instead of CE-Dice loss which was used for the oversampling and reweighing experiments. The reason for this is that Group DRO uses losses from individual samples to calculate the average loss for the groups. However, Dice loss is calculated using global statistics of the true positives, false positives and false negatives for a group or batch. It is non-additive as the numerator and denominator will change if the calculation is performed on a per-sample basis rather than for a group or batch. For Dice loss, the average group loss and global loss are different, which causes instability in training.
4.6 Training Using Combinations of Mitigation Methods
We also combined the mitigation methods into pairs to test whether combinations of methods would improve performance. This results in three additional methods: oversampling + Group DRO, reweighing + Group DRO, and oversampling + reweighing. These methods were applied in the same way as above. CE loss was used for experiments with Group DRO and CE-Dice loss was used for oversampling + reweighing.
@@ 4.7 Training Using Cropped Images
Following the findings of Lee et al. (2025a), which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images. The images were cropped around the heart using a bounding box defined based on a segmentation mask. All images were cropped to the same size, i.e. the size of the largest heart in the training set plus a buffer of 5 pixels in both the $x$ and $y$ directions.
When using cropped images, we evaluated two different approaches. First, we used the ground truth segmentations to calculate the cropping region at both training and inference time. Note that such a technique could not be used when the model is deployed since ground truth segmentations would not be available at inference time. Therefore, we evaluate this method to establish an upper bound on the performance of the cropping-based mitigation approach.
Second, we developed a ‘cascaded’ cropping-based approach, in which the cropping region was estimated using an initial nnU-Net-based segmentation. This first nnU-Net model was trained using the full images and its output was used to estimate the cropping region. The resulting cropped image was then used as input to a second nnU-Net model which was trained using cropped images. The croppings used on the training data in this second nnU-Net were based the ground truth segmentations, which are available at training time. This approach is illustrated in Fig. 1.
Figure 1: ‘Cascaded’ approach to bias mitigation based on using cropped images. Images (a) are first segmented using a full-image nnU-Net (b) to produce a segmentation (c). This segmentation is then used to crop the images (d) and used in another nnU-Net model trained using ground truth segmentation-based cropped images (e) to produce the final segmentations (f).
4.8 Evaluation
For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) and Hausdorff distance (HD) for subjects in the internal and external validation test sets. Performance was quantified using the median and inter-quartile range (IQR) of the DSC/HD. Unless otherwise stated, we report the values of these metrics averaged over the LVBP, LVM and RVBP.
Fairness metrics are also reported. The fairness gap (FG), represents the difference in median DSC between the protected groups (i.e. ethnicities) and is given by $FG=D_{White}-D_{Black}$ where $D$ is the median DSC. The skewed error ratio (SER), as defined in Puyol-Antón et al. (2021), is given by $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$. This measures the ratio of the errors between the median DSCs for the worst-performing and best-performing protected groups. For both fairness metrics a low value represents less bias and a more fair model will have a FG closer to 0 and a SER closer to 1.